The number of teachers at a local high school is 360 less than one-third of the number of students. If the total number of students and teachers at the school is 1200, how many of them are (a) teachers and (b) students?
Answer by fractalier(6550)

Call the number of students and teachers, s and t. Then we write
t = (1/3)s - 360 and
t + s = 1200 (or t = 1200 - s)
Now substitute and get
1200 - s = (1/3)s - 360
Add 360 and add s to both sides...
1560 = (4/3)s
Now multiply by 3/4 and get
s = (3/4)(1560) = 1170 students
which means there are
t = 30 teachers
(Not a healthy ratio...)